# Foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems and their applications for multi-strength intelligent classifiers

### v3

@@ 2.2 Inclusion relationships of hesitant fuzzy sets
The following statements hold for $A,B,C\in HF(U)$,
(1) If $A\subset_{p}B$ and $B\subset_{p}C$, then $A\subset_{p}C$.
(2) If $A\subset_{a}B$ and $B\subset_{a}C$, then $A\subset_{a}C$.
(3) If $A\subset_{m}B$ and $B\subset_{m}C$, then $A\subset_{m}C$.
(4) If $A\subset_{s}B$ and $B\subset_{s}C$, then $A\subset_{s}C$.
(5) If $A\subset_{t}B$ and $B\subset_{t}C$, then $A\subset_{t}C$.
(6) If $A\subset_{n}B$ and $B\subset_{n}C$, then $A\subset_{n}C$.
@@
Let $\bar{A}$ and $\bar{B}$ be two classical sets, we know that $\bar{A}\sqcap\bar{B}\sqsubset\bar{A}$, $\bar{A}\sqcap\bar{B}\sqsubset\bar{B}$, $\bar{A}\sqsubset\bar{A}\sqcup\bar{B}$, $\bar{B}\sqsubset\bar{A}\sqcup\bar{B}$, and $\bar{A}\sqcap\bar{B}\sqsubset\bar{A}\sqcup\bar{B}$. However, these rules are not necessarily true for two hesitant fuzzy sets $A$ and $B$ (shown as Propositions 2.11-2.18).
@@
The following statements hold for $A,B\in HF(U)$,
(1) $A\cap B\subset_{p}A$.
(2) $A\cap B\subset_{p}B$.
(3) $A\subset_{p}A\cup B$.
(4) $B\subset_{p}A\cup B$.
(5) $A\cap B\subset_{p}A\cup B$.
Proof (1) $min\{A^{+}(x),B^{+}(x)\}\leqslant A^{+}(x)$.
(2) $min\{A^{+}(x),B^{+}(x)\}\leqslant B^{+}(x)$.
(3) $A^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$.
(4) $B^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$.
(5) $min\{A^{+}(x),B^{+}(x)\}\leqslant max\{A^{+}(x),B^{+}(x)\}$$\blacksquare$
Proposition 2.12.
The following statements hold for $A,B\in HF(U)$,
(1) $A\cap B\subset_{a}A$.
(2) $A\cap B\subset_{a}B$.
(3) $A\subset_{a}A\cup B$.
(4) $B\subset_{a}A\cup B$.
(5) $A\cap B\subset_{a}A\cup B$.
Proof (1) $min\{A^{+}(x),B^{+}(x)\}\leqslant A^{+}(x)$ and $min\{A^{-}(x),B^{-}(x)\}\leqslant A^{-}(x)$.
(2) $min\{A^{+}(x),B^{+}(x)\}\leqslant B^{+}(x)$ and $min\{A^{-}(x),B^{-}(x)\}\leqslant B^{-}(x)$.
(3) $A^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$ and $A^{-}(x)\leqslant max\{A^{-}(x),B^{-}(x)\}$.
(4) $B^{+}(x)\leqslant max\{A^{+}(x),B^{+}(x)\}$ and $B^{-}(x)\leqslant max\{A^{-}(x),B^{-}(x)\}$.
(5) $min\{A^{+}(x),B^{+}(x)\}\leqslant max\{A^{+}(x),B^{+}(x)\}$ and $min\{A^{-}(x),B^{-}(x)\}\leqslant max\{A^{-}(x),B^{-}(x)\}$$\blacksquare$
@@
$\sum\limits_{h\in(A\cap B)(x)}h=\sum\limits_{h\in A(x)}h+\sum\limits_{h\in B(x)}h=k_{1}\bar{a}+k_{2}\bar{b}\leqslant max(\bar{a},\bar{b})(k_{1}+k_{2})$. $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})\leqslant max(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=max(\bar{a},\bar{b})$. Then, one of $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ holds.
@@
$A^{+}(x)\neq B^{+}(x)$, let $V=\{h:h\in A(x)\ or\ h\in B(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime}=min(V)$, then $h^{\prime}>min(A^{+}(x),B^{+}(x))=max\{h:h\in(A\cap B)(x)\}$.
Since $h^{\prime}>max\{h:h\in(A\cap B)(x)\}$ and $|(A\cap B)(x)|+|V|=k_{1}+k_{2}$, then $\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$. Since $h^{\prime}=min(V)$, then $(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$.
Then, we can obtain $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})\leqslant max(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=max(\bar{a},\bar{b})$. Then, one of $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ holds.
@@
$\sum\limits_{h\in(A\cup B)(x)}h=\sum\limits_{h\in A(x)}h+\sum\limits_{h\in B(x)}h=k_{1}\bar{a}+k_{2}\bar{b}\geqslant min(\bar{a},\bar{b})(k_{1}+k_{2})$.
$\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})\geqslant min(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=min(\bar{a},\bar{b})$. Then, one of $A(x)\subset_{m}(A\cup B)(x)$ and $B(x)\subset_{m}(A\cup B)(x)$ holds.
@@
$A^{-}(x)\neq B^{-}(x)$, let $V=\{h:h\in A(x)\ or\ h\in B(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=max(V)$. $h^{\prime}<max(A^{-}(x),B^{-}(x))=min\{h:h\in(A\cup B)(x)\}$.
Since $h^{\prime}=max(V)$ and $|(A\cup B)(x)|+|V|=k_{1}+k_{2}$, then $(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$. Since $h^{\prime}<min\{h:h\in(A\cup B)(x)\}$, then $\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$.
Based on the results above, then $mean[(A\cup B)(x)]=\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})\geqslant min(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=min(\bar{a},\bar{b})$. Then, one of $A(x)\subset_{m}(A\cup B)(x)$ and $B(x)\subset_{m}(A\cup B)(x)$ holds.
@@
In this case, $|(A\cap B)(x)|<k_{1}+k_{2}$ and $|(A\cup B)(x)|=k_{1}+k_{2}$. Let $V=\{h:h\in(A\cup B)(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime}=min(V)$. $h^{\prime}>min(A^{+}(x),B^{+}(x))=max\{h:h\in(A\cap B)(x)\}$.
Since $|(A\cap B)(x)|+|V|=|(A\cup B)(x)|=k_{1}+k_{2}$ and $h^{\prime}>max\{h:h\in(A\cap B)(x)\}$, then $\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$. Since $h^{\prime}=min(V)$, then $(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})$.
@@
In this case, $|(A\cap B)(x)|=k_{1}+k_{2}$ and $|(A\cup B)(x)|<k_{1}+k_{2}$. Let $V=\{h:h\in(A\cap B)(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=max(V)$. $h^{\prime}<max(A^{-}(x),B^{-}(x))=min\{h:h\in(A\cup B)(x)\}$.
Since $|(A\cup B)(x)|+|V|=|(A\cap B)(x)|=k_{1}+k_{2}$ and $h^{\prime}<min\{h:h\in(A\cup B)(x)\}$, then $\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$. Since $h^{\prime}=max(V)$, then $(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})$.
@@
Based on the condition $A^{-}(x)\neq B^{-}(x)$, let $V=\{h:h\in(A\cap B)(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=max(V)$, then $h^{\prime}<max(A^{-}(x),B^{-}(x))=min\{h:h\in(A\cup B)(x)\}$, and $|V|+|(A\cup B)(x)|=k_{1}+k_{2}$.
@@
Based on the condition $A^{+}(x)\neq B^{+}(x)$, let $W=\{h:h\in(A\cup B)(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime\prime}=min(W)$, then $h^{\prime\prime}>min(A^{+}(x),B^{+}(x))=max\{h:h\in(A\cap B)(x)\}$, and $|W|+|(A\cap B)(x)|=k_{1}+k_{2}$.
@@
For random two hesitant fuzzy sets $A$ and $B$, Example 2.14 shows that $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ do not hold simultaneously, $A(y)\subset_{m}(A\cup B)(y)$ and $B(y)\subset_{m}(A\cup B)(y)$ do not hold simultaneously.
$(A\cap B)(x)\subset_{m}B(x)$ and $(A\cap B)(x)\not\subset_{m}A(x)$. Furthermore, $(A\cap B)(z)\subset_{m}A(z)$ and $(A\cap B)(z)\not\subset_{m}B(z)$. It means that (1) of Proposition 2.13 cannot be written as that one of $A\cap B\subset_{m}A$ and $A\cap B\subset_{m}B$ holds.
@@
(1) At least one of $A(x)\subset_{s}(A\cup B)(x)$ and $A(x)\subset_{t}(A\cup B)(x)$ holds.
(2) At least one of $B(x)\subset_{s}(A\cup B)(x)$ and $B(x)\subset_{t}(A\cup B)(x)$ holds.
(3) At least one of $(A\cap B)(x)\subset_{s}(A\cup B)(x)$ and $(A\cap B)(x)\subset_{t}(A\cup B)(x)$ holds.
Proof (1) For each $x\in U$, if $A^{-}(x)\geqslant B^{-}(x)$, then $h\in(A\cup B)(x)$ for all $h\in A(x)$, i.e., $A(x)$ is a subsequence of $(A\cup B)(x)$. $A(x)\preccurlyeq A(x)$ and $|A(x)|\leqslant|(A\cup B)(x)|$, by Remark 2.8 (1) and (2), then one of $A(x)\subset_{s}(A\cup B)(x)$ and $A(x)\subset_{t}(A\cup B)(x)$ holds.
If $A^{-}(x)<B^{-}(x)$, let $K=A(x)\sqcap(A\cup B)(x)$, then $max(A(x)-K)<min(K)\leqslant max(K)<max((A\cup B)(x)-K)$, by Remark 2.8 (4), then one of $A(x)\subset_{s}(A\cup B)(x)$ and $A(x)\subset_{t}(A\cup B)(x)$ holds.
(2) It can be proved by the similar method of (1).
(3) Let $K=(A\cap B)(x)\sqcap(A\cup B)(x)$, then $max((A\cap B)(x)-K)<min(K)\leqslant max(K)<min((A\cup B)(x)-K)$, by Remark 2.8 (4), one of $(A\cap B)(x)\subset_{s}(A\cup B)(x)$ and $(A\cap B)(x)\subset_{t}(A\cup B)(x)$ holds.$\blacksquare$
Proposition ${}^{\prime}$ The following statements hold for $A,B\in HF(U)$,
(1) $A\subset_{sot}A\cup B$.
(2) $B\subset_{sot}A\cup B$.
(3) $A\cap B\subset_{sot}A\cup B$.
Proposition 2.16.
The following statements hold for $x\in U$ and $A,B,C\in HF(U)$,
@@
(6) For a random $C\in HF(U)$, if $A\subset_{s}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
(7) For a random $C\in HF(U)$, if $A\subset_{t}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
(8) For a random $C\in HF(U)$, if $A\subset_{n}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
@@
(5) $A\subset_{n}B$ implies $A\subset_{p}B$, by the result of (1), then (5) holds.
@@
Based on the illustrations of sub-case $C^{-}(x)\leqslant B^{-}(x)$, we have $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$, where $q=min\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.8 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
@@
Based on the illustrations of sub-case $B^{-}(x)<C^{-}(x)$, we have $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$, where $q=min\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.8 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.

### v4

@@ 2.3 Some propositions about the intersection and union of two hesitant fuzzy sets
The following statements hold for $A_{1},A_{2}\in HF(U)$,
(1) $A_{1}\cap A_{2}\subset_{p}A_{i}$ for $i=1,2$.
(2) $A_{i}\subset_{p}A_{1}\cup A_{2}$ for $i=1,2$.
(3) $A_{1}\cap A_{2}\subset_{p}A_{1}\cup A_{2}$.
@@
(3) $inf\{A_{1}^{+}(x),A_{2}^{+}(x)\}\leqslant sup\{A_{1}^{+}(x),A_{2}^{+}(x)\}$$\blacksquare$
@@
The following statements hold for $A_{1},A_{2}\in HF(U)$,
(1) $A_{1}\cap A_{2}\subset_{a}A_{i}$ for $i=1,2$.
(2) $A_{i}\subset_{a}A_{1}\cup A_{2}$ for $i=1,2$.
(3) $A_{1}\cap A_{2}\subset_{a}A_{1}\cup A_{2}$.
Proof (1) and (2) are obvious.
(3) $inf\{A_{1}^{+}(x),A_{2}^{+}(x)\}\leqslant sup\{A_{1}^{+}(x),A_{2}^{+}(x)\}$ and $inf\{A_{1}^{-}(x),A_{2}^{-}(x)\}\leqslant sup\{A_{1}^{-}(x),A_{2}^{-}(x)\}$$\blacksquare$
@@
$\sum\limits_{h\in(A\cap B)(x)}h=\sum\limits_{h\in A(x)}h+\sum\limits_{h\in B(x)}h=k_{1}\bar{a}+k_{2}\bar{b}\leqslant sup(\bar{a},\bar{b})(k_{1}+k_{2})$. $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})\leqslant sup(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=sup(\bar{a},\bar{b})$. Then, one of $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ holds.
@@
$A^{+}(x)\neq B^{+}(x)$, let $V=\{h:h\in A(x)\ or\ h\in B(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime}=inf(V)$, then $h^{\prime}>inf(A^{+}(x),B^{+}(x))=sup\{h:h\in(A\cap B)(x)\}$.
Since $h^{\prime}>sup\{h:h\in(A\cap B)(x)\}$ and $|(A\cap B)(x)|+|V|=k_{1}+k_{2}$, then $\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$. Since $h^{\prime}=inf(V)$, then $(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$.
Then, we can obtain $mean[(A\cap B)(x)]=\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})\leqslant sup(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=sup(\bar{a},\bar{b})$. Then, one of $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ holds.
@@
$\sum\limits_{h\in(A\cup B)(x)}h=\sum\limits_{h\in A(x)}h+\sum\limits_{h\in B(x)}h=k_{1}\bar{a}+k_{2}\bar{b}\geqslant inf(\bar{a},\bar{b})(k_{1}+k_{2})$.
$\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})\geqslant inf(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=inf(\bar{a},\bar{b})$. Then, one of $A(x)\subset_{m}(A\cup B)(x)$ and $B(x)\subset_{m}(A\cup B)(x)$ holds.
@@
$A^{-}(x)\neq B^{-}(x)$, let $V=\{h:h\in A(x)\ or\ h\in B(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=sup(V)$. $h^{\prime}<sup(A^{-}(x),B^{-}(x))=inf\{h:h\in(A\cup B)(x)\}$.
Since $h^{\prime}=sup(V)$ and $|(A\cup B)(x)|+|V|=k_{1}+k_{2}$, then $(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})$. Since $h^{\prime}<inf\{h:h\in(A\cup B)(x)\}$, then $\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$.
Based on the results above, then $mean[(A\cup B)(x)]=\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant(k_{1}\bar{a}+k_{2}\bar{b})/(k_{1}+k_{2})\geqslant inf(\bar{a},\bar{b})(k_{1}+k_{2})/(k_{1}+k_{2})=inf(\bar{a},\bar{b})$. Then, one of $A(x)\subset_{m}(A\cup B)(x)$ and $B(x)\subset_{m}(A\cup B)(x)$ holds.
@@
In this case, $|(A\cap B)(x)|<k_{1}+k_{2}$ and $|(A\cup B)(x)|=k_{1}+k_{2}$. Let $V=\{h:h\in(A\cup B)(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime}=inf(V)$. $h^{\prime}>inf(A^{+}(x),B^{+}(x))=sup\{h:h\in(A\cap B)(x)\}$.
Since $|(A\cap B)(x)|+|V|=|(A\cup B)(x)|=k_{1}+k_{2}$ and $h^{\prime}>sup\{h:h\in(A\cap B)(x)\}$, then $\sum\limits_{h\in(A\cap B)(x)}h/|(A\cap B)(x)|<(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$. Since $h^{\prime}=inf(V)$, then $(\sum\limits_{h\in(A\cap B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\leqslant\sum\limits_{h\in(A\cup B)(x)}h/(k_{1}+k_{2})$.
@@
In this case, $|(A\cap B)(x)|=k_{1}+k_{2}$ and $|(A\cup B)(x)|<k_{1}+k_{2}$. Let $V=\{h:h\in(A\cap B)(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=sup(V)$. $h^{\prime}<sup(A^{-}(x),B^{-}(x))=inf\{h:h\in(A\cup B)(x)\}$.
Since $|(A\cup B)(x)|+|V|=|(A\cap B)(x)|=k_{1}+k_{2}$ and $h^{\prime}<inf\{h:h\in(A\cup B)(x)\}$, then $\sum\limits_{h\in(A\cup B)(x)}h/|(A\cup B)(x)|>(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})$. Since $h^{\prime}=sup(V)$, then $(\sum\limits_{h\in(A\cup B)(x)}h+h^{\prime}|V|)/(k_{1}+k_{2})\geqslant\sum\limits_{h\in(A\cap B)(x)}h/(k_{1}+k_{2})$.
@@
Based on the condition $A^{-}(x)\neq B^{-}(x)$, let $V=\{h:h\in(A\cap B)(x),h\not\in(A\cup B)(x)\}$ and $h^{\prime}=sup(V)$, then $h^{\prime}<sup(A^{-}(x),B^{-}(x))=inf\{h:h\in(A\cup B)(x)\}$, and $|V|+|(A\cup B)(x)|=k_{1}+k_{2}$.
@@
Based on the condition $A^{+}(x)\neq B^{+}(x)$, let $W=\{h:h\in(A\cup B)(x),h\not\in(A\cap B)(x)\}$ and $h^{\prime\prime}=inf(W)$, then $h^{\prime\prime}>inf(A^{+}(x),B^{+}(x))=sup\{h:h\in(A\cap B)(x)\}$, and $|W|+|(A\cap B)(x)|=k_{1}+k_{2}$.
@@
For two random hesitant fuzzy sets $A$ and $B$, Example 2.16 shows that $(A\cap B)(x)\subset_{m}A(x)$ and $(A\cap B)(x)\subset_{m}B(x)$ do not hold simultaneously, $A(y)\subset_{m}(A\cup B)(y)$ and $B(y)\subset_{m}(A\cup B)(y)$ do not hold simultaneously.
$(A\cap B)(x)\subset_{m}B(x)$ and $(A\cap B)(x)\not\subset_{m}A(x)$. Furthermore, $(A\cap B)(z)\subset_{m}A(z)$ and $(A\cap B)(z)\not\subset_{m}B(z)$. It means that (1) of Proposition 2.15 cannot be written as that one of $A\cap B\subset_{m}A$ and $A\cap B\subset_{m}B$ holds. When we discuss the case ``$\subset_{m}"$ we need to note that
$\begin{cases}A_{1}\cap A_{2}\subset_{m}A_{i}\ \text{is uncertain for}\ i=1,2,\\ A_{i}\subset_{m}A_{1}\cup A_{2}\ \text{is uncertain for}\ i=1,2.\\ \end{cases}$
Proposition 2.17.
The following statements hold for $x\in U$ and $A_{1},A_{2}\in HF(U)$,
(1) At least one of $A_{i}(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $A_{i}(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds for $i=1,2$.
(2) At least one of $(A_{1}\cap A_{2})(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $(A_{1}\cap A_{2})(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds.
Proof (1) We prove the case that one of $A_{1}(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $A_{1}(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds in follows.
For each $x\in U$, if $A_{1}^{-}(x)\geqslant A_{2}^{-}(x)$, then $h\in(A_{1}\cup A_{2})(x)$ for all $h\in A_{1}(x)$, i.e., $A_{1}(x)$ is a subsequence of $(A_{1}\cup A_{2})(x)$. $A_{1}(x)\preccurlyeq A_{1}(x)$ and $|A_{1}(x)|\leqslant|(A_{1}\cup A_{2})(x)|$, by Remark 2.11 (1) and (2), then one of $A_{1}(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $A_{1}(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds.
If $A_{1}^{-}(x)<A_{2}^{-}(x)$, let $K=A_{1}(x)\sqcap(A_{1}\cup A_{2})(x)$, then $sup(A_{1}(x)-K)<inf(K)\leqslant sup(K)<sup((A_{1}\cup A_{2})(x)-K)$, by Remark 2.11 (4), then one of $A_{1}(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $A_{1}(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds.
With the same manner, we can prove the case that one of $A_{2}(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $A_{2}(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds.
(2) Let $K=(A_{1}\cap A_{2})(x)\sqcap(A_{1}\cup A_{2})(x)$, then $sup((A_{1}\cap A_{2})(x)-K)<inf(K)\leqslant sup(K)<inf((A_{1}\cup A_{2})(x)-K)$, by Remark 2.11 (4), one of $(A_{1}\cap A_{2})(x)\subset_{s}(A_{1}\cup A_{2})(x)$ and $(A_{1}\cap A_{2})(x)\subset_{t}(A_{1}\cup A_{2})(x)$ holds.$\blacksquare$
Proposition 2.17 The following statements hold for $A_{1},A_{2}\in HF(U)$,
(1) $A_{i}\subset_{sot}A_{1}\cup A_{2}$ for $i=1,2$.
(2) $A_{1}\cap A_{2}\subset_{sot}A_{1}\cup A_{2}$.
@@
(5) $A\subset_{n}B$ implies $A\subset_{p}B$, by the result of (1), then (5) holds.$\blacksquare$
Proposition 2.18 The following statements hold for $A,B,C\in HF(U)$,
(1) If $A\subset_{p}B$, then $A\subset_{sot}A\cap B$.
(2) If $A\subset_{a}B$, then $A\subset_{sot}A\cap B$.
(3) If $A\subset_{sot}B$, then $A\subset_{sot}A\cap B$.
(4) If $A\subset_{n}B$, then $A\subset_{sot}A\cap B$.
Example 2.19.
Proposition 2.20.
The following statements hold for $A,B\in HF(U)$,
(1) If $A\subset_{n}B$, $A\cap B\subset_{n}B$.
(2) If $A\subset_{n}B$, $A\subset_{n}A\cup B$.
(3) If $A\subset_{n}B$, $A\cap B\subset_{n}A\cup B$.
Example 2.21.
Let $U=\{x\}$, $A=\frac{\{0.1,0.3,0.5\}}{x}$, $B=\frac{\{0.2,0.4,0.6\}}{x}$.
2.4 Some propositions about a pair of hesitant fuzzy sets (one being a subset of the other) and a random additional hesitant fuzzy set
For three classical sets $\bar{A}$, $\bar{B}$ and $\bar{C}$, if $\bar{A}\sqsubset\bar{B}$, then $\bar{A}\sqsubset\bar{B}\sqcup\bar{C}$, $\bar{A}\sqcap\bar{C}\sqsubset\bar{B}\sqcap\bar{C}$ and $\bar{A}\sqcup\bar{C}\sqsubset\bar{B}\sqcup\bar{C}$. These rules in the cases of three hesitant fuzzy sets $A$, $B$ and $C$ are investigated, shown as Proposition 2.22 and 2.24.
Proposition 2.22.
The following statements hold for $x\in U$ and $A,B,C\in HF(U)$,
(1) If $A\subset_{p}B$, then $A\subset_{p}B\cup C$.
(2) If $A\subset_{a}B$, then $A\subset_{a}B\cup C$.
(3) If $A\subset_{s}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
(4) If $A\subset_{t}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
(5) If $A\subset_{n}B$, then at least one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
Proof (1) and (2) are obvious.
@@
Based on the illustrations of sub-case $C^{-}(x)\leqslant B^{-}(x)$, we have $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$, where $q=inf\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.11 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
@@
Based on the illustrations of sub-case $B^{-}(x)<C^{-}(x)$, we have $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$, where $q=inf\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.11 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.